OpenStudy (anonymous):

Help? An acute angle θ is in a right triangle with cos θ = 2/3. What is the value of sec θ?

OpenStudy (anonymous):

If the cos is 2/3, that means that adj. side is 2, and the hyp. is 3, right? That would mean the third side would have to be the sqrt of 13... Right?

OpenStudy (anonymous):

@surjithayer

OpenStudy (jdoe0001):

\(\bf recall\implies sec(\theta)=\cfrac{1}{cos(\theta)}\)

OpenStudy (anonymous):

I can't figure that out until I figure out the third side. Not a single of of the answer choices have sqrt of 13.. What am I doing wrong?

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{{\color{blue}{ \square }}}{{\color{red}{ \square }}}\qquad sec(\theta)=\cfrac{{\color{red}{ \square }}}{{\color{blue}{ \square }}}\)

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

Oh, wait. Cos is adj/hp, so I do 2^2 + b^2=3^2. 4+b=9 So the third side is sqrt of 5. My bad.

OpenStudy (anonymous):

Actually, if I had paid attention, I din't even need the third side... cos is the opposite of sec, so it's 3/2. Right?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thanks, hun. :P Sorry I'm slow sometimes. Haha. God bless you!